OpenStudy (anonymous):

Which ratio is equivalent to 4 : 13?

OpenStudy (muzzack):

8:26

OpenStudy (muzzack):

12:39

OpenStudy (muzzack):

any answer choices

OpenStudy (anonymous):

yes a. 8:17 b 16:52 c 12:33 d. 20:25

hero (hero):

Hint: \[\frac{4}{13} \times \frac{1}{1} = \frac{4}{13}\] \[\frac{4}{13} \times \frac{2}{2} = \frac{8}{26}\] \[\frac{4}{13} \times \frac{3}{3} = \frac{12}{39}\] See the pattern? Keep going until you match one of the answer choices.

OpenStudy (muzzack):

you multiply whatever number with 4 and to the other number basically

hero (hero):

I'm starting to become intolerant with the answer giving.

OpenStudy (muzzack):

sorry :(

hero (hero):

If anyone posts anything along the lines of "the answer is A, B, C, or D" then you are not helping anyone learn anything.

OpenStudy (muzzack):

im sosososososososososososososo sorry :(

OpenStudy (muzzack):

i dont mean it

OpenStudy (muzzack):

well what he said @Hero

hero (hero):

1. Avoid posting just the answer. Doing so does not help anyone learn anything. 2. Guide, Explain, Demonstrate, Detail 3. It is best to know and understand how to solve it yourself before attempting to help.

OpenStudy (muzzack):

this is a sequence A sequence is an ordered list. Like a set, it contains members. The number of ordered elements is called the length of the sequence.

OpenStudy (muzzack):

im so sorry mister @Hero

OpenStudy (muzzack):

let me expalin @samanthamaddenlove

OpenStudy (muzzack):

What is a Sequence? A Sequence is a set of things (usually numbers) that are in order. Sequence Infinite or Finite If the sequence goes on forever it is called an infinite sequence, otherwise it is a finite sequence

OpenStudy (muzzack):

in mathematics, a geometric series is a series with a constant ratio between successive terms. For example, the series is geometric, because each successive term can be obtained by multiplying the previous term by 1/2.